OpenStudy (anonymous):

When a product is raised to a power, it is equal to the product of each factor raised to that power. Show that this rule does not apply for a sum or difference to a power using the problem (5 – 3)2. You may indicate an exponent in your answer with ^. For example, 3x2 as 3x^2.

OpenStudy (anonymous):

Here are two ways of showing the value of the given expression:\[\begin{align*} (5-3)^2&=2^2\\ &=4 \end{align*}\] \[\begin{align*} (5-3)^2&=(5-3)(5-3)\\ &=5(5-3)-3(5-3)&\text{distributive property}\\ &=25-15-15+9&\text{dist prop}\\ &=4 \end{align*}\] Meanwhile, you have \[\begin{align*} (5-3)^2&\not=5^2-3^2\\ &=25-9\\ &=16\not=4 \end{align*}\]